The demand function P=4500-15Q2 and the supply function P = 5Q2 + 2500. Determine the following:



Consumer surplus.



Producer surplus.

1
Expert's answer
2022-03-07T11:14:54-0500

At equilibrium,


"4,500-15Q^2=5Q^2+2,500"


"4,500-2,500=5Q^2+15Q^2"


"2,000=20Q^2"


"\frac {2,000}{20} =Q^2"


"\sqrt100=\sqrt Q^2"


"Q=10"


Price is calculated below

"5(10)^2+2,500"


"500+2,500= 3,000"


Consumer surplus is calculated as


"\u222b_0^q\u200bd(Q)dQ\u2212P\\times Q"


"CS=∫_0 ^{10} 4,500−15Q^2dQ−{3,000\times 10}"


"=[4,500 Q\u2212\\frac{15Q^3}{3}\u200b]\u221230,000"


Since Q= 10;


"=[45,000\u22125,000]\u221230,000"


"=\$ 10,000"


producer surplus is calculated as;


"P\times Q−∫_0^q s(Q)dQ."


"PS={ 3,000\times 10}−∫_0^{10}5Q^2 +2,500 dQ"


"=30,000−(\frac{15Q^3}{3}+2,500Q)"


Since Q =10


"PS=30,000\u2212{5,000\u22122,500}"


"=\$ 0"